OpenStudy (anonymous):

Help please.. In Europe, gasoline efficiency is measured in km/L. If your car's gas mileage is 36.0mi/gal , how many liters of gasoline would you need to buy to complete a 142-km trip in Europe? Use the following conversions: 1km=0.6214mi and 1gal=3.78L.

OpenStudy (anonymous):

Solve : \[36\frac{ mi }{ gal } . \frac{ 1 km }{ 0.6214 mi } . \frac{ 1 gal }{ 3.78 L }\] then you get Km/L then you can divide 142Km by that answer and you'll get Liters of gasoline.

OpenStudy (vincent-lyon.fr):

Interesting conversion. For your information, what the question says is not true. In continental Europe, car efficiency is measured in liters of gasoline per 100km Hence an standard cars consumption would be around 5 to 8 L/100km